OpenStudy (anonymous):

Solve: −2(2x + 5) − 3 = −3(x − 1)

OpenStudy (anonymous):

distribute, add lke terms and isolate x on one side

OpenStudy (anonymous):

-4x + -10 - 3 = -3x +3 -4x - 13 = -3x +3 -4x +3x = 3+13 -x = 16 x = -16